OpenStudy (anonymous):

If 10 kilometers is equal to 6 miles, how many miles are in 25 kilometers? @ranga

OpenStudy (anonymous):

15 miles

OpenStudy (anonymous):

15miles

OpenStudy (anonymous):

25 km is 2 and a half lots of 6 miles

OpenStudy (anonymous):

It's 15 miles :)

OpenStudy (anonymous):

\[\frac{ 6 }{ 10 } = \frac{ x }{ 25 }\]

OpenStudy (lifeisadangerousgame):

6/10 = x/25 cross multiply 150 = 10x divide by 10 on both sides 15 = x 15 miles is your answer c:

OpenStudy (anonymous):

Oh okay , thaaanks guuyssss ! (: